Joe mows lawns in his neighborhood for extra money. Suppose the demand for lawn mowing in Joe’s neighborhood is: QD = 50 – 2P And that the supply of lawn mowing (Joe’s willingness to mow lawns) is: QS = -10 + P. The market-clearing price of lawn mowing is $_____20___ per lawn and the market-clearing quantity is ____10___ lawns. (Enter your responses rounded to two decimal places.) The corresponding consumer surplus is $______________.
Answer #1

Answer

the equilibrium is at Qd=Qs

50-2P=-10+P

3P=60

P=20

Q=50-2*20=10

The market-clearing price of lawn mowing is $20.00 per lawn and the market-clearing quantity is 10.00 lawns

===========

converting the demand curve to inverse demand curve to find the consumer reservation price which is the y-axis intercept of the curve and constant term in the inverse equation

Q=50-2P

2P=50-Q

P=25-0.5Q

Consumer surplus =0.5*(Consumer reservation price -P)*Q

=0.5*(25-20)*10

=$25

the consumer surplus is $25.00